OpenStudy (anonymous):

A square is inscribed in a circle. The area of the square is 64 square inches. How many square inches are in the area of the circle? Express your answer in terms of pi.

OpenStudy (anonymous):

is A square inside a circle?

myininaya (myininaya):

i got 128pi

myininaya (myininaya):

i will post how i got it

OpenStudy (anonymous):

\[A circle= 8\pi\]than: A square = 8^2=64 Dimeter = 8 Area of the circle =( pi*8) \[ A =8 \pi\]

OpenStudy (anonymous):

sorry I was wrong

myininaya (myininaya):

oops i made a mistake

myininaya (myininaya):

32pi

OpenStudy (anonymous):

A square = 8^8 =64\[\sqrt{4^{2}+4^{2}}= \sqrt{32} 32pi is correct answer

myininaya (myininaya):

nancy did you see my cool work on this problem i like this problem

OpenStudy (anonymous):

thank,good cluck your home work

myininaya (myininaya):

yes good luck dream

OpenStudy (anonymous):

Thank you both very much!!!